# Guaranteed Recovery of Unambiguous Clusters

### v1

@@ VI Experiments
We compare a modified version of Algorithm 1 for handling overlapping clusters to widely used algorithms for non-convex cluster recovery on a range of datasets. This modified version uses a variation of Algorithm 2 where each time a cluster is output, the points in the cluster are removed from $X$. This makes the check of whether a new cluster intersects with a previously output cluster inapplicable since only unclustered points remain in $X$ at the beginning of each greedy step. To increase speed, instead of finding the min-$A$ clustering exactly, the implementation approximates the min-$A$ clustering by progressively adjusting $A$ until a $K$-clustering is output by the modified version of Algorithm 2. To improve performance slightly, we use $N_{p}=2$ in Algorithm 3, regardless of the $N_{p}$ used for finding the initial partial clusters.
@@
We compare our algorithm to $K$-means, spectral clustering, HDBSCAN, OPTICS, and spectACl [13]. We refer to our algorithm as “MSE” which stands for minimal seed expansion, and implemented it in Python. We use the spectral clustering and $K$-means implementations from scikit-learn, where the number of clusters $K$ is specified by the user. The affinity matrix in spectral clustering is formed using each point’s $K_{n}$ nearest neighbors, where $K_{n}$ is by the user. The implementation of $K$-means uses the “greedy $K$-means++” algorithm [17, 18]. We use the HDBSCAN implementation from the Python HDBSCAN clustering library, and we run it using the default cluster selection criteria named Excess of Mass (eom). The default setting for HDBSCAN sets the $N_{p}$ equal to $M+1$ where $M$ is the minimum possible cluster size set by the user. We refer to this default version as “HDBSCAN.” We refer to the version of HDBSCAN where we $N_{p}$ and $M+1$ are not tied as “HDBSCAN (2).” We use the OPTICS implementation from scikit-learn, where $N_{p}$, the minimum cluster size $M$, and the cluster selection parameter $X_{i}$ are set by the user. Neither HDBSCAN nor OPTICS uses the number of clusters $K$. We use the spectACl implementation available on the TU Dortmund website from the authors (https://sfb876.tu-dortmund.de/spectacl/index.html), and use the default setting where $\epsilon$ is set by the user.
@@
The Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI) are the most common measures of similarity between an estimated clustering and a ground truth clustering, and for each algorithm, we report both for each dataset tested. The algorithms’ performance is compared on the real world benchmark datasets whose properties are in Table I. All of these datasets are available on the UCI server [19]. “Cancer” refers to the Breast Cancer Wisconsin (Diagnostic) dataset. “Digits” refers to the test set of the Optical Recognition of Handwritten Digits dataset. “Letters” refers to the test set of the Letter Recognition dataset. “MNIST” refers to the test set of the MNIST dataset. For the MNIST dataset, we used t-SNE to reduce the dimensionality to two [20]. We also compare the same algorithms for 5 artificial datasets given in Figure 3 which were used in the example titled “Comparing different clustering algorithms on toy datasets” on the scikit-learn website [21].
@@
TABLE I: benchmark datasets and their properties
The results for benchmark datasets are reported in Table II, and results for artificial datasets are in Table IV. For MSE, we did not optimize $N_{p}$, $M$ and $D$ exhaustively. Instead, we set $N_{p}=3$, picked $M$ smaller than the true minimum cluster size for each dataset, and used the $D$ value from the set $\{1.5,2,10\}$ that gave the best ARI.
@@
For spectral clustering, we tried all $K_{n}$ in the range $\{1,2,...,20\}$ and reported the clustering with the best ARI. For HDBSCAN, we tried all $N_{p}$ in the range $\{2,3,...,20\}$ and reported the clustering with best ARI. For HDBSCAN (2) and OPTICS, we used the same $M$ and $N_{p}$ we used for our algorithm. For OPTICS, we tried all $X_{i}$ values in the set $\{0,0.05,0.1,...,0.95\}$ and reported the clustering with best ARI. For spectACl, we tried all $\epsilon$ values in the range $\{0,0.1,0.2,...,100\}$. For some datasets where the optimal $\epsilon$ was less than 20, we tried all $\epsilon$ values in the range $\{0,0.1,0.2,...,20\}$.
@@
For spectral clustering, in order of dataset appearance in the Table II, the chosen $K_{n}$ values are $4,6,18,3,5,4,19,11$. For HDBSCAN, in order of appearance in the Table II, the chosen $N_{p}$ values are $3,20,6,3,6,4,3,14$. For MSE, HDBSCAN (2), and OPTICS, $N_{p}$ was set to 3 for all datasets and $M$ was set to 35 for Iris, 30 for Wine, 50 for Seeds, 8 for Glass, 100 for Cancer and Digits, 70 for Letters, and 600 for MNIST. For OPTICS, in order of appearance in the Table II, the chosen $X_{i}$ values are $0,0.25,0,0.2,0.05,0,0,0$. For spectACl, in order of appearance in the Table II, the chosen $\epsilon$ values are $3.3,44.3,2.0,2.3,77.2,33.2,8.9,9.9$. For our algorithm, $D$ was set to 20 for Iris, Wine, Glass, and Cancer, 2 for Seeds, Letters and MNIST, and 1.5 for Digits.
@@
The results for artificial datasets in Figure 3 are reported in Table IV. For spectral clustering, in order of appearance in Table IV, the chosen $K_{n}$ values are $5,5,18,6,15$. For HDBSCAN, in order of appearance in Table IV, the chosen $N_{p}$ values are $3,6,9,9,14$. For MSE, HDBSCAN (2), and OPTICS, $N_{p}$ was set to 3 for all datasets and $M$ was set to 60 for all datasets in Table IV. For OPTICS, in order of appearance in the Table IV, the chosen $X_{i}$ values are $0.15,0.1,0.3,0.2,0.15$. For spectACl, in order of appearance in Table IV, the chosen $\epsilon$ values are $0.2,0.2,1.2,0.8,1.3$. For MSE, $D$ was set to 10 for Varied Variance Blobs, and 2 for all other datasets in Table IV.
@@
TABLE III: artificial datasets and their properties
Figure 3: Artificial datasets used in Table IV .
@@
TABLE IV: $|\hat{C}|$ is the number of clusters output by the algorithm. For HDBSCAN and OPTICS, the set of noise points counts as a cluster.
Observe that with just a lower bound on the minimum cluster size ($M$), and some knowledge of the maximum difference in maximum density among different clusters ($D)$, our algorithm gives as good or better ARI than the other algorithms on all benchmark datasets except the Glass dataset. On the Glass dataset, HDBSCAN gives the best ARI, but outputs an incorrect number of clusters. The NMI values obtained by our algorithm are also competitive. MSE performs strongly on the artificial datasets as well. Compared to the other algorithms, MSE performs well more consistently across the datasets tested. As an example, in the case of the Digits dataset, it is clear in Figure 4 that our algorithm does a significantly better job of recovering the Digits clustering than the other algorithms. While our implementation of MSE was not optimized for speed, MSE clustered each dataset in roughly three minutes or less on a laptop computer with 32 threads and 64 GB of RAM.
For MSE, $D$ (and similarly $M$ and $N_{p}$) can be optimized by running MSE for various values of $D$ and using the value that gives the clustering with the best DBCV score [22] (or other cluster quality score).
@@ VII Conclusion
In this paper we propose an information-theoretic characterization of when a $K$-clustering is ambiguous, and design an algorithm called Minimal Seed Expansion (MSE) that provably recovers the clustering whenever it is unambiguous. This characterization formalizes the situation when two high density regions within a cluster are separable enough that they look more like two distinct clusters than two truly distinct clusters in the clustering. We then implement and test a version of MSE that is modified to effectively handle overlapping clusters, and observe that it displays improved performance on many datasets with little variation in parameter selection across datasets tested. This improvement is in comparison to widely used algorithms for non-convex cluster recovery whose parameters are optimized independently for every dataset tested. MSE also performs well more consistently than all other algorithms tested in these experiments.

### v3

@@ VI Experiments
We compare a modified version of Algorithm 1 for handling overlapping clusters to widely used algorithms for non-convex cluster recovery on a range of datasets. This modified version uses a variation of Algorithm 2 where each time a cluster is output, the points in the cluster are removed from $X$. This makes the check of whether a new cluster intersects with a previously output cluster inapplicable since only unclustered points remain in $X$ at the beginning of each greedy step. To increase speed, instead of finding the minimum $A$ that outputs a $K$-clustering, the implementation approximates this value of $A$ by progressively adjusting $A$ until a $K$-clustering is output by the modified version of Algorithm 2. To improve performance slightly, we use $N_{p}=2$ in Algorithm 3, regardless of the $N_{p}$ used for finding the initial partial clusters. We implemented this modified version of Algorithm 1 in Python, and refer to it as “MSE” in this section. Treating the dimensionality of the data as a constant, this implementation runs in $O(t|X|^{2})$ time where $t$ is the number of $A$ values tried. The code for this implementation is available at: https://github.com/kmazooji/Minimal-Seed-Expansion.
We compare our algorithm to $K$-means, spectral clustering, HDBSCAN, OPTICS, and SpectACl [13]. Treating the dimensionality of the data as a constant and letting $t$ be the number of iterations of $K$-means, $K$-means runs in $O(tK|X|)$ time, spectral clustering with k-nearest neighbors affinity runs in $O(K|X|^{2}+tK|X|)$ time, HDBSCAN runs in $O(|X|^{2})$ time, OPTICS runs in $O(|X|^{2})$ time, and SpectACl runs in $O(d|X|^{2}+tK|X|)$ time where $d$ is the embedding dimension.
@@
We use the spectral clustering and $K$-means implementations from scikit-learn, where the number of clusters $K$ is specified by the user. The affinity matrix in spectral clustering is formed using each point’s $K_{n}$ nearest neighbors, where $K_{n}$ is by the user. The implementation of $K$-means uses the “greedy $K$-means++” algorithm [17, 18]. We use the HDBSCAN implementation from the Python HDBSCAN clustering library, and we run it using the default cluster selection criteria named Excess of Mass (eom). The default setting for HDBSCAN sets $N_{p}$ equal to $M+1$ where $M$ is the minimum possible cluster size set by the user. We refer to this default version as “HDBSCAN.” We refer to the version of HDBSCAN where $N_{p}$ and $M+1$ are not tied as “HDBSCAN (2).” We use the OPTICS implementation from scikit-learn, where $N_{p}$, the minimum cluster size $M$, and the cluster selection parameter $X_{i}$ are set by the user. Neither HDBSCAN nor OPTICS uses the number of clusters $K$. SpectACl uses knowledge of $K$ and accepts a parameter $\epsilon$. We use the implementation available on the TU Dortmund website from the authors: https://sfb876.tu-dortmund.de/spectacl/index.html.
@@
The Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI) are the most common measures of similarity between an estimated clustering and a ground truth clustering. For each algorithm, we report both measures for each dataset tested. The algorithms’ performance is compared on the real world benchmark datasets whose properties are in Table I. All of these datasets are available on the UCI server [19]. “Cancer” refers to the Breast Cancer Wisconsin (Diagnostic) dataset. “Digits” refers to the test set of the Optical Recognition of Handwritten Digits dataset. “Letters” refers to the test set of the Letter Recognition dataset. “MNIST” refers to the test set of the MNIST dataset. For the MNIST dataset, we used t-SNE to reduce the dimensionality to two [20]. We also compare the same algorithms for 5 artificial datasets given in Figure 4 which were used in the example titled “Comparing different clustering algorithms on toy datasets” on the scikit-learn website [21]. The properties of the datasets are given in Table IV.
TABLE I: benchmark datasets and their properties
@@
The first set of results for benchmark datasets are reported in Table II. In these experiments, for MSE, we did not optimize $N_{p}$, $M$ and $D$ exhaustively. Instead, we set $N_{p}=3$, picked $M$ smaller than the true minimum cluster size for each dataset to give a competitive ARI, and used the $D$ value from the set $\{1.5,2,20\}$ that gave the best ARI.
@@
TABLE III: $|\hat{C}|$ is the number of clusters output by the algorithm.
dataset algorithm ARI NMI $|\hat{C}|$
Iris MSE 0.886 0.871 3
MSE (auto) 0.835 0.833 3
Wine MSE 0.439 0.430 3
MSE (auto) 0.359 0.420 3
Seeds MSE 0.725 0.682 3
MSE (auto) 0.725 0.682 3
Glass MSE 0.232 0.379 7
MSE (auto) 0.232 0.379 7
Cancer MSE 0.743 0.628 2
MSE (auto) 0.694 0.595 2
Digits MSE 0.864 0.898 10
MSE (auto) 0.819 0.879 10
Letters MSE 0.193 0.447 26
MSE (auto) 0.193 0.447 26
@@
For the other algorithms tested, we used grid search over the parameters to maximize ARI. For spectral clustering, we tried all $K_{n}$ in the range $\{1,2,...,20\}$ and reported the clustering with the best ARI. For HDBSCAN, we tried all $N_{p}$ in the range $\{2,3,...,20\}$ and reported the clustering with best ARI. For HDBSCAN (2), we used the same $M$ and $N_{p}$ we used for MSE. For OPTICS, we used the same $N_{p}$ value that we used for MSE, and tried all $X_{i}$ values in the set $\{0,0.05,0.1,...,0.95\}$ and all $M$ values in the set $\{\delta\cdot|X|/K:\delta\in\Delta\}$ where $\Delta=\{0.05,0.10,0.15,...,1\}$, and reported the clustering with best ARI. For MNIST, we tested $\Delta=\{0.1,0.2,0.3,...,1\}$ for OPTICS to save time. For SpectACl, we tried all $\epsilon$ values in the range $\{0,0.1,0.2,...,100\}$. For some large datasets where the optimal $\epsilon$ was clearly less than 20, we tried all $\epsilon$ values in the range $\{0,0.1,0.2,...,20\}$.
@@
For MSE, and HDBSCAN (2), $N_{p}$ was set to 3 for all datasets and $M$ was set to 35 for Iris, 30 for Wine, 50 for Seeds, 3 for Glass, 100 for Cancer and Digits, 70 for Letters, and 600 for MNIST. For MSE, $D$ was set to 20 for Iris, Wine, Glass, and Cancer, 2 for Seeds, Letters and MNIST, and 1.5 for Digits. For spectral clustering, in order of dataset appearance in Table II, the chosen $K_{n}$ values are $4,6,18,3,5,4,19,11$. For HDBSCAN, in order of appearance in Table II, the chosen $N_{p}$ values are $3,20,6,3,6,4,3,14$. For OPTICS, in order of appearance in Table II, the chosen $M$ values are $38,47,56,20,57,162,54,800$, and the chosen $X_{i}$ values are $0,$ $0,$ $0,$ $0.15,$ $0.05,$ $0,$ $0,$ $0.05$. For SpectACl, in order of appearance in Table II, the chosen $\epsilon$ values are $3.3,$ $44.3,$ $2.0,$ $2.3,$ $77.2,$ $33.2,$ $8.9,$ $9.9$.
The second set of results for benchmark datasets is reported in Table III. MSE (auto) refers to MSE where $M$ values and $D$ values are optimized to give the best Calinski-Harabasz score, which is an internal clustering metric, meaning that it does not use the ground truth labels to asses clustering quality. For all of these experiments, we set $N_{p}=3$. We chose the $M$ from $\{\delta\cdot|X|/K\;:\;\delta\in\Delta\}$ where $\Delta=[0.025,0.05,0.075,...,0.975]$ and chose $D$ from $[1.5,2,20]$. We chose from these values of $M$ because we know the number of clusters $K$, and we choose from these values of $D$ because we have observed at least one of these values to work well on a wide range of datasets. In Table III, we also reported the statistics for the MSE clusterings from Table II for comparison. We observe that in general, the decrease in clustering quality is not large on these datasets. The average decrease in ARI is 0.032, and the average decrease in NMI is 0.014. The average percent decrease in ARI is 5.1% and the average percent decrease in NMI is 2.0%. We did not test on MNIST due to the slower speed of the algorithm on this dataset. Note that we only tested this approach for this one internal clustering quality measure, and the results may improve if another internal clustering quality measure (e.g. DBCV score [22]) is used to optimize the clustering. Additionally, $N_{p}$ can also be optimized using this approach.
@@
The results for artificial datasets in Figures 4 are reported in Table V. For MSE and HDBSCAN (2), $N_{p}$ was set to 3 for all datasets, and $M$ was set to 60 for all datasets in Table V. For MSE, $D$ was set to 10 for Varied Variance Blobs, and 2 for all other datasets in Table V. For spectral clustering, in order of appearance in Table V, the chosen $K_{n}$ values are $5,5,18,6,15$. For HDBSCAN, in order of appearance in Table V, the chosen $N_{p}$ values are $3,6,9,9,14$. For OPTICS, $N_{p}$ was set to 3, and in order of appearance in Table V, the chosen $M$ values are $225,225,133,167,167$, and the chosen $X_{i}$ values are $0.0,0.0,0.05,0.0,0.1$ For SpectACl, in order of appearance in Table V, the chosen $\epsilon$ values are $0.2,0.2,1.2,0.8,1.3$.
TABLE IV: artificial datasets and their properties
@@
Figure 4: Artificial datasets used in Table V .
TABLE V: $|\hat{C}|$ is the number of clusters output by the algorithm. For HDBSCAN and OPTICS, the set of noise points counts as a cluster.
@@
Observe that with a lower bound on the minimum cluster size ($M$), and some knowledge of the maximum difference in maximum density among different clusters ($D)$, MSE gives better ARI than the other algorithms on all benchmark datasets except Glass, and gives competitive ARI for all artificial datasets. On the Glass dataset, OPTICS gives the best ARI, but outputs an incorrect number of clusters. The NMI values obtained by our algorithm are also competitive. Compared to the other algorithms, MSE performs well more consistently across the datasets tested, despite the fact that the parameters were not optimized in an exhaustive manner. As an example, in the case of the Digits dataset, it is clear in Figure 5 that our algorithm does a significantly better job of recovering the Digits clustering than the other algorithms. While our implementation of MSE was not optimized for speed, MSE clustered the MNIST dataset in three minutes on a laptop computer with 32 threads and 64 GB of RAM. MSE ran much faster on the other datasets.
@@ VII Conclusion
In this work, we propose an information-theoretic characterization of when a $K$-clustering is ambiguous, and design an algorithm called Minimal Seed Expansion (MSE) that provably recovers the clustering whenever it is unambiguous. This characterization formalizes the situation when two high density regions within a cluster are separable enough that they look more like two distinct clusters than two truly distinct clusters in the clustering. We then implement and test a version of MSE that is modified to effectively handle overlapping clusters, and observe that it displays improved performance on many datasets without its parameters exhaustively optimized for each dataset tested. This improvement is in comparison to widely used algorithms for non-convex cluster recovery whose parameters are optimized using grid search independently for every dataset tested. MSE also performs well more consistently than all other algorithms tested in these experiments. We also optimized the parameters for MSE using grid search to maximize the Calinski-Harabasz score (an internal clustering quality measure), and observed little decrease in ARI and NMI compared to the case where parameters were chosen to give competitive ARI. This suggests that MSE can also be used effectively without manual parameter tuning.
VIII Acknowledgments
The work of K.M. and I.S. was supported in part by the National Science Foundation CAREER Award under Grant CCF-2046991. The authors thank the anonymous reviewers of ISIT 2025 for their helpful comments.